OpenStudy (annidix2):

Write the equation in standard form. 8−5y=−2x

OpenStudy (anonymous):

-5y=-2x-8 divide by -5 on each side to isolate y y=2/5x+8/5

OpenStudy (triciaal):

standard form Ax + By = C

OpenStudy (anonymous):

first find x

OpenStudy (calculusxy):

standard form: \(ax + by = c\) in the equation \(8 - 5y = -2x\) you have the \(by\) and the \(c\) terms on one side, but you have the \(ax\) on the other side of the equal sign. you need to move them so that the equation looks like \(ax + by = c\)

OpenStudy (triciaal):

@mtran1 that is correct for the slope-intercept form y = m x + b

OpenStudy (calculusxy):

@mtran1 there is a difference between the slope intercept form (\(y = mx + b\)) and the standard form (\(ax + by = c\))

OpenStudy (anonymous):

you all are to solve for x and y not finding the slope formular

OpenStudy (annidix2):

i tryed y=2/5x+8/5 but it said it was rong

OpenStudy (triciaal):

it is wrong because that is the slope-intercept form and the question asked for standard form

OpenStudy (triciaal):

again, also expressed by @calculusxy the standard form is Ax + By = C rearrange the original to match this format and you will get it right

OpenStudy (triciaal):

@Daniel56k you cannot solve a linear equation with 2 unknowns

OpenStudy (anonymous):

okay